OpenStudy (anonymous):

The sum of an infinite number of terms of GP is 4 and the sum of their cubes is 192. Find the series.

OpenStudy (anonymous):

\[a+aq+aq^2+...=4\]\[a^3+a^3q^3+a^3q^6+...=192\]

OpenStudy (anonymous):

so first one gives\[\frac{a}{1-q}=4\]and the second gives\[\frac{a^3}{1-q^3}=192\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

if u want check...\[q=-\frac{1}{2}\]